The length of each side of an equilateral triangle is 5 centimeters more than the length of each side of a square. The perimeter of the two figures are equal. Find the lengths of the of the square and of the triangle.
I really don't know where to start all I know is that you have to define the variable and write a formula and use it to solve the problem. Please and thank you !

Answer by ankor@dixie-net.com(22740)

The length of each side of an equilateral triangle is 5 centimeters more than the length of each side of a square. The perimeter of the two figures are equal. Find the lengths of the of the square and of the triangle.
:
Let x = side of the square
:
It says,"The length of each side of an equilateral triangle is 5 centimeters more than the length of each side of a square", therefore:
:
(x+5) = side of the triangle.
:
"The perimeter of the two figures are equal. "
4x = 3(x+5)
:
4x = 3x + 15
:
4x - 3x = 15
:
x = 15
:
:
Is this true, find the perimeter of each; (15 + 5 = 20, side of the triangle)
4(15) = 60
3(20) = 60